# Critical Evaluation of Studies Alleging Evidence for Technosignatures in the POSS1-E Photographic Plates

### v1

@@ III.2 Principal uses and additional properties
Set $R$: This set is the tiny residuum that survived the Solano et al. (2022) filter pipeline, which was designed to remove all known celestial objects and digitization-related artifacts from $S$. As the most highly filtered subset of $A$, we subjected $R$ to closer scrutiny by manually examining cropped POSS1-E images and Pan-STARRS images of the same field (published online by the authors [^3] See http://svocats.cab.inta-csic.es/vanish-possi/; in what follows, we supply ID numbers for the filenames of downloadable images on this site). We reviewed 10% of the SPFs in $R$ (i.e., 540 images), finding 12 that appear to be of objects with exceptionally high proper motion (i.e., exceeding the $5^{\prime\prime}$ search radius) that appear in the Pan-STARRS frame (1043, 1051, 1059, 1130, 1243, 1271, 1277, 1345, 1374, 1376, 1409, 1513), and 11 that appear to form part of an obvious artifact (1092, 1248, 1255, 1460, 1179, 1223, 1259, 1301, 1396, 1473, 1507). Examples of these are shown in Figure 2. We therefore estimate that between 4% and 5% (i.e., $\sim 25/540$) of the objects in $R$ are catalog stars or obvious artifacts, which suggests that the overwhelming majority are roughly symmetrical SPFs without a catalog match.
@@
Set $V$: Given that set $R$ is currently the most vetted dataset of unidentified SPFs, it is not clear why Villarroel et al. (2025a) and Bruehl and Villarroel (2025) resampled $P$ $(N=126,412)$ to create a new set $V$ ($N=107,185$). Set $P$ includes 116,607 SPFs within 5′′ of an optical catalog object and 410 SPFs within 5′′ of an infrared catalog object, which means $V$ is known to comprise at least 91% SPFs that have not yet been distinguished from catalog objects. Furthermore, $V$ has not passed through the Solano et al. (2022) scan artifact removal step. We estimate up to $\sim 40$% of the objects in $V$ would be classified by Solano et al. (2022) as scan artifacts, based on the fraction identified and removed to obtain $R$ as reported in that study (see Table 1).
@@
POSS1-E plates: There is no mention in Solano et al. (2022) about the removal of SPFs or plates imaged in the southern celestial hemisphere, which leaves unexplained why the southern hemisphere plates were not sampled or, if they were sampled, why these SPFs were removed and at what stage in the construction of sets $S,P,W,R$ and $V$.
@@ IV Spatial distribution
Before proceeding, we define the phrase “spatially uniform-random distribution” or simply “uniform-random” as referring to the spatial distribution resulting from a Poisson point process: e.g., the spatial distribution characteristic of star fields. We define “nonrandom” as any spatial distribution of features that, in principle, can be distinguished from “uniform-random” using a statistical test.
@@ IV.1 Variation in counts per plate
We examine the variation of counts per plate as part of assessing assumptions implicit in the analyses of Villarroel et al. (2025a) and Bruehl and Villarroel (2025) about what factors supposedly influence the distribution of SPF counts in time and space, such as the terrestrial shadow and nuclear tests. Primary sources of variation in counts per plate should be characterzied, modeled, and subtracted before second-order effects are considered. In particular, we might expect the number of SPFs per plate to vary for two primary reasons. First, if the SPFs represent a population of stars not removed by prior filtering, then we expect (i) a roughly uniform-random distribution with an overall increase in number density with decreasing distance from the Galactic plane; and (ii) slight variations associated with variations in the astronomical viewing conditions (i.e., more stars are detected under favorable atmospheric conditions with less turbulence blur). Alternatively, if the SPFs represent plate artifacts, the distribution will depend on the processes that tend to create these features, and upon how the impact of these processes varies from plate to plate. This may in turn depend on how the plates were manufactured, exposed, developed, stored, and handled, and how often they have been accessed. Finally, we should consider that the digital scanning process may also introduce artifacts, such as by introducing images of dust or debris in the scanner or on the plate surface. Solano et al. (2022) removed digitization artifacts only after removing over 117,000 SPFs from set $P$ (NN=126,412) when constructing set $R$ ($N=5,399$), whereas the removal of digitization artifacts was neither explicitly described nor mentioned in Villarroel et al. (2025a) or Bruehl and Villarroel (2025) as part of the sampling of $V$ ($N=107,875$) from $P$.
@@ IV.2 Distribution with respect to plate boundaries
Figure 4: Scatter plots of SPFs from $W$ on nine plates that exhibit distinctive patterns. Red dots signify SPFs that could not be unambiguously assigned to the plate in question because of overlap with the neighboring plate’s field of view (and hence confined to the edges). These patterns are not visible in the dataset $M$ of SPFs that appear in both E and O plates (i.e., likely celestial objects), suggesting that they are unrelated to nonuniform sensitivity of the instrumentation or photographic medium. The geometric patterns and relationship to plate boundaries strongly suggest (i) a physical process that has templated the spatial distribution of SPFs, such as a surface that has come into contact with the plate, or a region that has been preferentially exposed to processes that tend to degrade the medium over time; or (ii) a spatial filter applied when processing the data. We have plotted SPFs from $W$ because (i) it is publicly accessible; (ii) it is large enough for these patterns to be visible, and (iii) because it is a dense, uniform-random sample and majority subset of $S$, the parent set of set $V$ analyzed by Bruehl and Villarroel (2025) and Villarroel et al. (2025a). Each plate is labeled with the plate ID, the number of SPFs on the plate, and the value of the Evans-Clark ratio $\rho$, which is smallest for the most overdispersed (clustered) distributions.
@@
Figure 5: Histograms of the Evans-Clark ratio $\rho$ calculated for each plate, for SPFs in sets $M$ and $W$. As expected for a uniform-random distribution of points on a finite domain, $\rho$ is typically slightly above 1 for set $M$, comprising SPFs of likely celestial origin because they appear in both the POSS1-E (red band) and POSS1-O plates (blue band). The distribution of SPFs in $W$ is dominantly overdispersed (clustered), with $\rho<1$ for most plates. See Figure 4 for scatter plots of SPFs on individual plates with significant overdispersion, labeled with the estimated $\rho$.
@@
Figure 6: Radial density profiles for the $W$ (solid), $R$ (dashed), and $M$ (dotted) datasets. Densities were calculated in circular bins of $0.25^{\circ}$ with radii denoting the angular distance from plate centers. All distributions were normalized for easy visual comparison of their shapes. The $W$ and $R$ distributions show an increasing density of SPFs away from plate centers, in qualitative agreement with the experimental results of Greiner et al. (1987), who found that the number density of endemic plate emulsion defects increases with distance from plate centers. By contrast, $M$ exhibits a slight and gradual decline with radius, indicating a mostly uniform number density combined with the effect of vignetting, as expected for a star field.
@@ IV.3 Templating and spatial filtering
We hypothesize that the distribution of SPFs in some regions of some plates are “templated” by contact with objects or surfaces, or by preferential exposure to a process that tends to degrade the photographic medium. The fact that SPFs occur within regions with clear boundaries, or else avoid regions with clear boundaries, suggests that contact or exposure-based templating may play an important role in the production of SPFs on many other plates, where the contact (or exposed) surface mostly or completely covered them. An example zone of avoidance is visible in the upper-left corner in part G of Figure 4. Based on inspection of multiple plates, the vertical stripes in parts E and F span the interval in right ascension from roughly $103^{\circ}$ to $107^{\circ}$ and cross plate boundaries; a second stripe from roughly $86^{\circ}$ to $98^{\circ}$ also crosses plate boundaries. Both of these features are further discussed in Section IV.4.
We emphasize the importance of answering the following questions. First, why do SPFs in some cases appear to reside in clusters or regions of higher number density with clear boundaries, and why do they sometimes appear to avoid regions with clear boundaries? Second, why do these overdense regions and empty regions in some cases appear related to the plate footprint or the celestial coordinate system (e.g., Plate 090R in part A of Figure 4)? Third, what does this imply about the production of SPFs in general? These questions should be answered as part of the essential data validation process.
@@ IV.4 Distribution across the sky
We call attention to this feature of the on-sky distribution because analyses in Villarroel et al. (2025a) and Bruehl and Villarroel (2025) implicitly assume that $V$ is completely and uniformly sampled (i.e., no part of the north celestial hemisphere was omitted when SPFs were selected) and faithfully records the distribution of SPFs across the sky. This assumption is essential for one of the key claims in Villarroel et al. (2025a), which reports a deficit of SPFs from set $V^{\prime}$ in the Earth’s shadow, as discussed in the next section. This is also significant for interpreting the background rate of SPFs made in the same paper. An incomplete or nonuniform sampling would affect correlation statistics between SPF occurrence and any other phenomena, such as nuclear tests Bruehl and Villarroel (2025).
There are marked similarities and differences evident from comparing the spatial distributions of $W$ and $R$ (see parts B and C of Figure 8), which shows the effect of removing additional catalog objects and scan artifacts from $S$ by Solano et al. (2022). Across the western celestial hemisphere, SPFs in the set $R$ ($N=5,399)$ tend to hug the plane of the Galaxy, suggesting that some fraction of these objects may be uncataloged stars, including uncatalogued dwarf stars that were transiently flaring during the POSS1-E exposures. There is also a marked depletion north of the celestial equator in the western hemisphere when compared to the eastern hemisphere in both $W$ and $R$ (white dashed outline in Figure 8). In the case of $R$, SPFs in the western hemisphere above the Galactic plane have been almost totally erased. Before using $R$ to search for images of unidentified objects, the origin of these patterns in the spatial distribution should be investigated and understood.
@@ IV.5 Clusters of aligned SPFs
At the start of Section 4 in Villarroel et al. (2025a), the authors write “We base our analysis on the catalog of 298,165 short duration transients presented in Solano et al. (2022), detected in red POSS1 plates with typical exposure times of 45–50 minutes.” This implies that the search for aligned SPFs made use of $S$ ($N=298,165$), which according to Solano et al. (2022) is 96.8% composed of features that reside within $5^{\prime\prime}$ of a catalog star (including what we have called set $W$). Indeed, as shown in our Table 2, we find that 8 of the 24 SPFs belonging to candidate linear clusters in Table 3 of Villarroel et al. (2025a) reside within 2′′ of an object in $W$ (a subset of $S$), suggesting that these features represent one and the same object. Villarroel et al. (2025a) recomputed the astrometric positions of SPFs in Table 3 (Villarroel et al., 2025a) with respect to positions originally reported for SPFs in $W$ from Solano et al. (2022), which likely accounts for the $2^{\prime\prime}$ positional differences we see here. We have confirmed that 8 of 24 SPFs in Table 3 of Villarroel et al. (2025a) are located within 5′′ of a NeoWISE catalog object position (a 9th is $\sim 5.1^{\prime\prime}$ away). These members of set $W$ were removed from $S$ by Solano et al. (2022) when constructing $R$ because they were not confidently distinguished from NeoWISE objects. For the same reason, Villarroel et al. (2025a) also removed $W$ from $S$ to construct the set $V$ used in other analyses. No reason is given in Villarroel et al. (2025a) for reverting to set $S$ to search for SPF alignments representing satellite glints.
@@
We manually examined the SuperCosmos and DSS images of all SPFs in Table 3 of Villarroel et al. (2025a) and show two comparisons in Figure 9. We have noted a marked qualitative asymmetry of some of these SPFs, indicating the need for a quantitative morphological analysis. (A morphological analysis is also recommended by the presence of the anomalous shapes of some SPFs in $R$ as shown in Figure 2.) In Figure 8 of Villarroel et al. (2025b), the lowermost member of the cluster of three SPFs (red arrows in part B of our Figure 9) is described as “slightly dubious in shape”. In our high-magnification view, this is clearly an asymmetrical feature. As an important reminder, the SPFs in $S$ have not been aggressively filtered for a high degree of radial symmetry. These images also highlight two defects of the kind that can be removed by scan comparison, in this case appearing on the SuperCosmos scan and not the DSS scan (yellow arrows). These artifacts are also clearly identified in Figure 8 of Villarroel et al. (2025a). Parts C and D of our Figure 9 show another SPF from an aligned cluster (object 3 in cluster 5 of Table 3 in Villarroel et al. (2025a)) that also exhibits clear elongation, although is possibly a pair of overlapping sources. An additional star-like SPF artifact appears in the lower right (yellow arrow).
Figure 9: Magnified images of SPFs belonging to candidate linear clusters detailed in Table 3 of Villarroel et al. (2025a) (red arrows) and SPF scan artifacts present in SuperCosmos scans (B and D) and absent from DSS scans (A and C) of the same field (yellow arrows). The noticeable departure from circular symmetry in the case of multiple objects marked with red arrows suggests that $S$ should be aggressively filtered to remove features without an extremely high degree of radial symmetry, as expected for the point source glints hypothesized in Villarroel et al. (2025a), in order to avoid likely artifacts.

### v2

@@ III.2 Principal uses and limitations
Set $R$: This set is the tiny residuum that survived the Solano et al. (2022) filter pipeline, which was designed to remove all known celestial objects and digitization-related artifacts from $S$. As the most highly filtered subset of $A$, we subjected $R$ to closer scrutiny by manually examining cropped POSS1-E images and Pan-STARRS images of the same field; in what follows, we supply ID numbers for the filenames of downloadable images on this site). We reviewed 10% of the SPFs in $R$ (i.e., 540 images), finding 12 that appear to be of objects with exceptionally high proper motion (i.e., exceeding the $5^{\prime\prime}$ search radius) that appear in the Pan-STARRS frame (1043, 1051, 1059, 1130, 1243, 1271, 1277, 1345, 1374, 1376, 1409, 1513), and 11 that appear to form part of an obvious artifact (1092, 1248, 1255, 1460, 1179, 1223, 1259, 1301, 1396, 1473, 1507). Examples of these are shown in Figure 2. We therefore estimate that between 4% and 5% (i.e., $\sim 25/540$) of the objects in $R$ are catalog stars or obvious artifacts, which suggests that the overwhelming majority are roughly symmetrical SPFs without a catalog match.
@@
Set $V$: Given that set $R$ is currently the most vetted dataset of unidentified SPFs, it is not clear why Villarroel et al. (2025a) and Bruehl and Villarroel (2025) resampled $P$ $(N=126,412)$ to create a new set $V$ ($N=107,185$). Set $P$ includes 116,607 SPFs within 5′′ of an optical catalog object and 410 SPFs within 5′′ of an infrared catalog object, which means $V$ is known to comprise at least 91% SPFs that have not yet been distinguished from catalog objects. Furthermore, $V$ has not passed through the Solano et al. (2022) scan artifact removal step. We previously estimated up to $\sim 40$% of the objects in $V$ would be classified by Solano et al. (2022) as scan artifacts, based on the fraction identified and removed to obtain $R$ as reported in that study (see Table 1).
@@
POSS1-E plate scans: There is no mention in Solano et al. (2022) about the removal of SPFs or plates imaged in the southern celestial hemisphere, which leaves unexplained why the southern hemisphere plates were not sampled or, if they were sampled, why these SPFs were removed and at what stage in the construction of sets $S,P,W,R$ and $V$.
@@ IV Spatial distribution
Before proceeding, we define the phrase “spatially uniform-random distribution” or simply “uniform-random” as referring to the spatial distribution resulting from a Poisson point process: e.g., the spatial distribution characteristic of star fields. We define “nonrandom” as any spatial distribution of features that, in principle, can be distinguished from “uniform-random” using a statistic like the Clark-Evans ratio (Clark and Evans, 1954).
@@ IV.1 Variation in counts per plate
We examine the variation of counts per plate as part of assessing assumptions implicit in the analyses of Villarroel et al. (2025a) and Bruehl and Villarroel (2025) about what factors supposedly influence the distribution of SPF counts in time and space, such as the terrestrial shadow and nuclear tests. Primary sources of variation in counts per plate should be characterized, modeled, and subtracted before second-order effects are considered. In particular, we might expect the number of SPFs per plate to vary for two primary reasons. First, if the SPFs represent a population of stars not removed by prior filtering, then we expect (i) a roughly uniform-random distribution with an overall increase in number density with decreasing distance from the Galactic plane; and (ii) slight variations associated with variations in the astronomical viewing conditions (i.e., more stars are detected under favorable atmospheric conditions with less turbulence blur). Alternatively, if the SPFs represent plate artifacts, the distribution will depend on the processes that tend to create these features, and upon how the impact of these processes varies from plate to plate. This may in turn depend on how the plates were manufactured, exposed, developed, stored, and handled, and how often they have been accessed. Finally, we should consider that the digital scanning process may also introduce artifacts, such as by introducing images of dust or debris in the scanner or on the plate surface. Solano et al. (2022) removed digitization artifacts only after removing over 117,000 SPFs from set $P$ (NN=126,412) when constructing set $R$ ($N=5,399$), whereas the removal of digitization artifacts was neither explicitly described nor mentioned in Villarroel et al. (2025a) or Bruehl and Villarroel (2025) as part of the sampling of $V$ ($N=107,875$) from $P$.
@@ IV.2 Distribution with respect to plate boundaries
Figure 4: Scatter plots of SPFs from $W$ on nine plates that exhibit distinctive patterns. Red dots signify SPFs that could not be assigned to a single plate because of overlap with the neighboring plate’s field of view. These patterns are not visible in the dataset $M$ of SPFs that appear in both E and O plates (i.e., likely celestial objects), suggesting that they are unrelated to nonuniform sensitivity of the instrumentation or photographic medium. The geometric patterns and relationship to plate boundaries strongly suggest (i) a physical process that has templated the spatial distribution of SPFs, such as a surface that has come into contact with the plate, or a region that has been preferentially exposed to processes that tend to degrade the medium over time; or (ii) a spatial filter applied when processing the data. We have plotted SPFs from $W$ because (i) it is publicly accessible; (ii) it is large enough for these patterns to be visible, and (iii) because it is a dense, uniform-random sample and majority subset of $S$, the parent set of set $V$ analyzed by Bruehl and Villarroel (2025) and Villarroel et al. (2025a). Each plate is labeled with the plate ID, the number of SPFs on the plate, and the value of the Evans-Clark ratio $\rho$, which is smallest for the most overdispersed (clustered) distributions.
@@
Figure 5: Histograms of the Evans-Clark ratio $\rho$ calculated for each plate, for SPFs in sets $M$ and $W$. As expected for a uniform-random distribution of points on a finite domain, $\rho$ is typically slightly above 1 for set $M$, comprising SPFs of likely celestial origin because they appear in both the POSS1-E (red band) and POSS1-O plates (blue band). The distribution of SPFs in $W$ is dominantly overdispersed (clustered), with $\rho<1$ for most plates. See Figure 4 for examples of scatter plots of SPFs on individual plates with significant overdispersion, labeled with the estimated $\rho$.
@@
Figure 6: Radial density profiles for the $W$ (solid), $R$ (dashed), and $M$ (dotted) datasets. Densities were calculated in circular bins of $0.25^{\circ}$ with radii denoting the angular distance from plate centers. All distributions were normalized for easy visual comparison of their shapes. The $W$ and $R$ distributions show an increasing density of SPFs away from plate centers, in qualitative agreement with the experimental results of Greiner et al. (1987), who found that the number density of endemic plate emulsion defects increases with distance from plate centers. By contrast, $M$ exhibits a slight and gradual decline with increasing radius, indicating a mostly uniform number density combined with the effect of vignetting, as expected for a star field.
@@ IV.3 Templating and spatial filtering
The fact that some SPFs occur within regions with clear boundaries, or else avoid regions with clear boundaries, suggests that contact or exposure-based templating may play a role in the production of SPFs more generally. An example zone of avoidance is visible in the upper-left corner in part G of Figure 4. Based on inspection of multiple plates, the vertical stripes in parts E and F span the interval in right ascension from roughly $103^{\circ}$ to $107^{\circ}$ and cross plate boundaries; a second stripe from roughly $86^{\circ}$ to $98^{\circ}$ also crosses plate boundaries. Both of these features are further discussed in Section IV.4.
We emphasize the importance of answering the following questions. First, why do SPFs in some cases appear to reside in clusters or regions of higher number density with clear boundaries, and why do they sometimes appear to avoid regions with clear boundaries? Second, why do these overdense regions and empty regions in some cases appear related to the plate footprint or the celestial coordinate system (e.g., Plate 090R in part A of Figure 4)? Third, what does this imply about the production of SPFs in these POSS1-derived datasets? These questions should be answered as part of the essential data validation process.
@@ IV.4 Distribution across the sky
We call attention to this feature of the on-sky distribution because analyses in Villarroel et al. (2025a) and Bruehl and Villarroel (2025) implicitly assume that $V$ is completely and uniformly sampled (i.e., no part of the north celestial hemisphere was omitted when SPFs were selected) and faithfully records the distribution of SPFs across the sky. This assumption is essential for one of the key claims in Villarroel et al. (2025a), which reports a deficit of SPFs from set $V^{\prime}$ in the Earth’s shadow, as discussed in the next section. This is also significant for interpreting the background rate of SPFs estimated in the same paper. An incomplete or nonuniform sampling would affect correlation statistics between SPF occurrence and any other phenomena, such as nuclear tests in Bruehl and Villarroel (2025).
There are marked similarities and differences evident from comparing the spatial distributions of $W$ and $R$ (see parts B and C of Figure 8), which shows the effect of removing additional catalog objects and scan artifacts from $S$ by Solano et al. (2022). Across the western celestial hemisphere, SPFs in the set $R$ ($N=5,399)$ tend to hug the plane of the Galaxy, suggesting that some fraction of these objects may be uncataloged stars, including uncatalogued dwarf stars that were transiently flaring during the POSS1-E exposures. Alternatively, the high number density of stars in this region result in blended objects with inaccurate centroid positions that frustrate the identification and removal of catalog objects. There is also a marked depletion north of the celestial equator in the western hemisphere when compared to the eastern hemisphere in both $W$ and $R$ (white dashed outline in Figure 8). In the case of $R$, SPFs in the western hemisphere above the Galactic plane have been almost totally erased. Before using $R$ to search for images of unidentified objects, the origin of these patterns in the spatial distribution should be investigated and understood.
@@ IV.5 Clusters of aligned SPFs
The start of Section 4 in Villarroel et al. (2025a) states that: “We base our analysis on the catalog of 298,165 short duration transients presented in Solano et al. (2022), detected in red POSS1 plates with typical exposure times of 45–50 minutes.” This implies that the search for aligned SPFs made use of $S$ ($N=298,165$), which according to Solano et al. (2022) is 96.8% composed of features that reside within $5^{\prime\prime}$ of a catalog star (including what we have called set $W$). Indeed, as shown in our Table 2, we find that 8 of the 24 SPFs belonging to candidate linear clusters in Table 3 of Villarroel et al. (2025a) reside within 2′′ of an object in $W$ (a subset of $S$), suggesting that these features represent one and the same object. Villarroel et al. (2025a) recomputed the astrometric positions of SPFs in Table 3 (Villarroel et al., 2025a) with respect to positions originally reported for SPFs in $W$ from Solano et al. (2022), which likely accounts for the $2^{\prime\prime}$ positional differences we see here. We have confirmed that 8 of 24 SPFs in Table 3 of Villarroel et al. (2025a) are located within 5′′ of a NeoWISE catalog object position (a 9th is $\sim 5.1^{\prime\prime}$ away). These members of set $W$ were removed from $S$ by Solano et al. (2022) when constructing $R$ because they were not confidently distinguished from NeoWISE objects. For the same reason, Villarroel et al. (2025a) also removed $W$ from $S$ to construct the set $V$ used in other analyses. No reason is given in Villarroel et al. (2025a) for reverting to set $S$ to search for SPF alignments representing satellite glints.
@@
We manually examined the SuperCosmos and DSS images of all SPFs in Table 3 of Villarroel et al. (2025a) and show two comparisons in Figure 9. We have noted a marked qualitative asymmetry of some of these SPFs, indicating the need for a quantitative morphological analysis in addition to the symmetry-based filtering already applied in Solano et al. (2022). (A morphological analysis is also recommended by the presence of the anomalous shapes of some SPFs in $R$ as shown in Figure 2.) In Figure 8 of Villarroel et al. (2025b), the lowermost member of the cluster of three SPFs (red arrows in part B of our Figure 9) is described as “slightly dubious in shape”. In our high-magnification view, this is clearly an asymmetrical feature and not representative of a star-like optical transient. These images also highlight two defects of the kind that can be removed by scan comparison, in this case appearing on the SuperCosmos scan and not the DSS scan (yellow arrows). These artifacts are also clearly identified in Figure 8 of Villarroel et al. (2025a). Parts C and D of our Figure 9 show another SPF from an aligned cluster (object 3 in cluster 5 of Table 3 in Villarroel et al. (2025a)) that also exhibits clear elongation, although is possibly a pair of overlapping sources.
Figure 9: Magnified images of SPFs belonging to candidate linear clusters detailed in Table 3 of Villarroel et al. (2025a) (red arrows) and SPF scan artifacts present in SuperCosmos scans (B and D) and absent from DSS scans (A and C) of the same field (yellow arrows). Parts A and B represent “candidate 3” and Parts C and D represent “candidate 5”. The noticeable departure from circular symmetry in the case of multiple objects marked with red arrows suggests that $S$ should be aggressively filtered to remove features without an extremely high degree of radial symmetry, as expected for the point source glints hypothesized in Villarroel et al. (2025a), in order to avoid likely artifacts.